2. c) A capillary tube with an inside radius of 638 nm can support a 100 mm column of liquid that has a density of 820 kgm3. The observed contact angle is 12.5°. Find the surface tension of the liquid for h= 20 cm.


1
Expert's answer
2021-12-06T09:46:40-0500

As we know,


"h=\frac{2\sigma\cos\theta}{\rho gr},\\\space\\
\sigma=\frac{\rho grh}{2\cos\theta}=5.25·10^{-4}\text{ N/m}."